OpenStudy (anonymous):

Find the zeros of f(x)= x(x-2)(x-1)

OpenStudy (anonymous):

set each factor equal to zero and solve

OpenStudy (anonymous):

x=2 x=1?

OpenStudy (anonymous):

you can pretty much do it in your head \(x=0\) or \(x-2=0\iff x=2\) or \(x-1=0\iff x=11\)

OpenStudy (anonymous):

oops i meant \(x=1\)

OpenStudy (anonymous):

don't forget \(x=0\)

OpenStudy (anonymous):

okay! Is that really it?

OpenStudy (anonymous):

don't forget \(x\) itself is a factor of \( f(x)= x(x-2)(x-1)\)

OpenStudy (anonymous):

that's the zero's?

OpenStudy (anonymous):

yeah,this is pretty much of a soft ball question isn't it?

OpenStudy (anonymous):

that is it zeros are \(\{0,1,2\}\) finis

OpenStudy (anonymous):

I guess so ! well thanks so much, I pulled a 4 A.M, and I just woke up and still working, she hasn't put in grades yet(: